OpenStudy (anonymous):

What is the solution set of |x – 4| + 7 = 4? –4 or 4 –3 or 3 No solution 1 or 7

OpenStudy (anonymous):

1

OpenStudy (nurali):

|x – 4| + 7 = 4 |x – 4|=4-7 |x – 4|=-3 x-4=-3 or x-4=3 x=1 or x=7

OpenStudy (anonymous):

wich one is it?

OpenStudy (nurali):

Last one.

OpenStudy (anonymous):

thank you I thought it was that one to wanted to make sure

OpenStudy (anonymous):

1 add 4 to -3 and the answer is 1